Economic state

Probability

T-Bill

Alta Indus.

Repo Men

American Foam

Market Port.(index)

Recession

0.1

8.00%

-22.0%

28.0%

10.0%

-13.0%

Below Average

0.2

8.00%

-2.0%

14.7%

-10.0%

1.0%

Average

0.4

8.00%

20.0%

0.0%

7.0%

15.0%

Above Average

0.2

8.00%

35.0%

-10.0%

45.0%

29.0%

Boom

0.1

8.00%

50.0%

-20.0%

30.0%

43.0

Barney Smith Investment Advisors recently issued estimates for the state of the economy and the rate of return on each state of the economy. Alta Industries, Inc. is an electronics firm; Repo Men Inc. collects past due debts; and American Foam manufactures mattresses and various other foam products. Barney Smith also maintains an "index fund" which owns a market-weighted fraction of all publicly traded stocks; you can invest in that fund and thus obtain average stock market results. Given the situation as described, answer the following questions.

a. Calculate the expected rate of return on each alternative.

b. Calculate the standard deviation of returns on each alternative.

c. Calculate the coefficient of variation on each alternative.

d. Calculate the beta on each alternative.

e. Do the SD, CV, and beta produce the same risk ranking? Why or why not?

f. Suppose you create a two-stock portfolio by investing $50,000 in Alta Industries and $50,000 in Repo Men. Calculate the expected return, standard deviation, coefficient of variation, and beta for this portfolio. How does the risk of this two-stock portfolio compare with the risk of the individual
Answer #1

Question a:

Expected rate of return of each alternative

1. Expected rate of return of T bill = 8% (since the return is 8% in all cases)

2. Expected return of Alta Indus. = 0.1*-22%+0.2*-2%+0.4*20%+0.2*35%+0.1*50% = 17.40%

3. Expected return of Repo Men = 0.1*28% +0.2*14%+0.4*0 +0.2*-10% +0.1*-20% = 1.6%

4. Expected return of American Foam = 0.1*10% +0.2*-10% +0.4*7%+0.2*45% +0.1*30% = 13.80%

5 Expected return of Market Portfolio = 0.1*-13% +0.2*1% +0.4*15% +0.2*29% +0.1*43% = 15.00%

Note: We have answered one full question with all the sub-parts (alternatives). Please note that only one full question can be answered at a time. Please post each question separately for experts to answer